### v3

[…]
Chenyu Huang1, Zhengyang Tang2,3, Dongdong Ge4, Shixi Hu5
[…]
2 The Chinese University of Hong Kong, Shenzhen (CUHK-Shenzhen)
3 Shenzhen Research Institute of Big Data
4 Shanghai Jiao Tong University
5 Cardinal Operations
6 Columbia University
[…]
rj2556@columbia.edu
{wangbenyou,wangzizhuo}@cuhk.edu.cn xin.zheng@duke.edu
[…]
Optimization modeling and solving play a critical role in the application of Operations Research (OR) tools to address real-world problems, yet they pose challenges and require extensive expertise from OR experts. With the advent of large language models (LLMs), new opportunities have emerged to streamline and automate these tasks. However, current research predominantly relies on closed-source LLMs such as GPT-4, along with extensive prompt engineering techniques. This reliance stems from the scarcity of high-quality training datasets for optimization modeling, resulting in elevated costs, prolonged processing times, and privacy concerns. To address these challenges, our work is the first to propose a viable path for training open-source LLMs that are capable of optimization modeling as well as developing and executing solver codes, eventually leading to a superior ability for automating optimization modeling and solving. Particularly, we introduce a semi-automated data synthesis framework designed for optimization modeling issues, named OR-Instruct. This framework merges the training data requirements of large models with the unique characteristics of optimization modeling problems, and allows for customizable enhancements tailored to specific scenarios or modeling types. To evaluate the performance of our proposed framework, we present the IndustryOR benchmark, the inaugural industrial standard for evaluating LLMs in solving practical OR problems. Utilizing data synthesized through OR-Instruct, we train various open-source LLMs with a capacity of 7 billion parameters (dubbed ORLMs). The resulting model demonstrates significantly enhanced optimization modeling capabilities, achieving state-of-the-art performance across the NL4OPT, MAMO, and IndustryOR benchmarks. Our code and data are available at https://github.com/Cardinal-Operations/ORLM.
[…]
Automated optimization modeling, large language model, synthetic data
1 Introduction
Optimization models have been a critical analytical tool for business decision-making, aiding companies in making optimal choices under various complex decision environments. With the fast development of computing capability (including both the development of algorithm and hardware) and the growing amount of data, optimization models are playing more and more significant roles in the operations of large corporations. For instance, optimization techniques used by JD.com enable them to handle 10 times the normal volume of orders during peak sales seasons, helping the company decrease its fulfillment expense ratio to a world-leading level of 6.5%, which has resulted in hundreds of millions of dollars in savings (Qin et al. 2022). Baosteel, one of the largest steel company in China, utilized integer optimization for production planning at its main plant in Shanghai, leading to a significant reduction in carbon monoxide emissions — over 500,000 tons annually — as reported by INFORMS (2013). The U.S. Census Bureau was recognized as a finalist for the 2022 Franz Edelman Award for their innovative use of operations research and analytics in reengineering field operations for the 2020 U.S. Census. This project utilizes advanced analytics, optimization, and machine learning techniques to automate and optimize the scheduling, workload assignments, and management of field data collection, a process previously done manually (Adams et al. 2023). Despite the huge amount of success stories, the sensitivity of optimization models to environmental changes necessitates frequent adjustments in dynamically evolving business contexts. This, combined with the high level of expertise and substantial labor costs required for optimization modeling, represents a significant bottleneck that has hindered the full potential to be unlocked for optimization techniques.
The advent of large language models (LLMs) gives an emerging hope to the above challenge. As soon as its advent, these pre-trained LLMs have exhibited remarkable capabilities in rapid knowledge comprehension and cross-disciplinary learning, demonstrating remarkable success and transformative potential across various fields (Huang et al. 2024a). For instance, in the medical sector, the Hua Tuo large model significantly aids in diagnostics and treatment processes (Wang et al. 2023b). In the realm of software development, Codex (Zaremba et al. 2021) excels in numerous programming languages, specifically designed to enhance coding tasks. Additionally, these models are effectively utilized in automated perceptual analysis (Li et al. 2024), providing legal advice (Cui et al. 2023), offering stock market trading strategies (Chen et al. 2022b), and improving recommendation systems (Xu et al. 2024a). In more specialized scientific areas, such as mathematical derivations, LLMs also show substantial potential for advanced reasoning. For example, Xu et al. (2024b) develop a pipeline that incorporates self-critique mechanisms to refine the capabilities of LLMs in mathematical derivation. Furthermore, Trinh et al. (2024) train Alphaproof with hundreds of billions of synthetic data used to tackle complex olympiad geometry problems. Similarly, MathScale (Tang et al. 2024) offers a straightforward and scalable method to generate high-quality data for mathematical reasoning, establishing a foundation for further training of large language models in mathematical reasoning capabilities. Xing et al. (2024) and Ahmed and Choudhury (2024) both emphasize the specific challenges in translating linguistic descriptions into mathematical formulations. Their studies highlight the superiority of advanced LLMs like GPT-4 in handling complex optimization problems and the potential need for specialized datasets and finetuning approaches to further enhance performance.
In the field of optimization modeling, a few studies have emerged in recent years attempting to automate the modeling and solving of optimization problems under natural language representations using pre-trained language models (PLMs) (Liu et al. 2019, Lewis et al. 2019). For example, NL4OPT (Ramamonjison et al. 2023) decomposes this task into multiple steps, where semantic entities are first recognized upon which mathematical models are formulated subsequently. However, this approach often results in accumulated errors in the pipeline and has limited generalization capabilities due to the relatively-small parameter scale. As LLMs become more advanced like the GPT-4 and Claude-3.5, researchers have begun to generate mathematical models by directly prompting these models. To extend to a complete solution that includes a program using mature solvers, prompting engineering methods such as Chain-of-Experts (Xiao et al. 2023), OptiMUS (AhmadiTeshnizi et al. 2024) and OptiGuide (Li et al. 2023a) employ multi-agent collaboration based on proprietary LLMs. These agents collaborate using complex reasoning chains to refine both the mathematical models and programs. Nonetheless, utilizing online APIs for advanced large models offers no assurances of privacy protection, which is prohibitive in some sensitive industry applications. Moreover, these approaches rely heavily on the capabilities of the foundational LLMs itself. As illustrated in Figure 1, even a basic optimization modeling problem presents challenges for the state-of-the-art foundational large language model, GPT-4. It demonstrates deficiencies in comprehending and transforming logical constraints, which imposes a limitation on methods dependent on these general closed-source LLMs.
Figure 1: An example is presented to illustrate the failure of advanced large models to accurately respond to seemingly straightforward optimization modeling questions. The responses highlighted in reds contain errors, specifically the introduction of non-linear terms and misunderstandings of logical relationships. This response was generated in June 2024, originating from the web-based GPT-4.
In summary, existing large language models (LLMs) for optimization modeling exhibit the following deficiencies:
  1. 1.
    Limited modeling capabilities. Contemporary advanced LLMs frequently exhibit errors in interpreting logical constraints or inadvertently incorporate nonlinear elements, complicating problem solution.
  2. 2.
    Insufficient quality of optimization modeling training data. The quality and scale of data are directly proportional to the capability of the model, as described by the Scaling Law (Kaplan et al. 2020). However, a shortage in high-quality optimization modeling training data is often a critical challenge, obstructing the advancement of LLMs’ modeling ability.
  3. 3.
    Data privacy concerns. Dependence on APIs for closed-source LLMs may lead to data breaches which hampers the broader adoption and development of LLM technologies in the industrial sector.
  4. 4.
    Test sets are relatively homogeneous. The efficacy of large language models in automatic modeling is predominantly assessed using the dataset provided by the NL4OPT competition (Ramamonjison et al. 2023). This dataset primarily focuses on simple linear programming tasks that exhibit relatively low complexity, narrow scope, and few types compared to real-world applications.
To fill this gap, this paper introduces, for the first time, a new path for training an open-source large language model specifically designed for modeling and solving optimization problems. To address the core of the above challenges, we present a semi-automated framework for the synthesis of high-quality optimization modeling data, aimed at progressively enhancing the specialized modeling capabilities of LLMs. Our main contributions can be summarized as follows:
  1. 1.
    To ensure the effectiveness, robustness, and real-world applicability of our model, we have identified four crucial requirements that the training dataset must fulfill, as informed by academic research like Alzubaidi et al. (2023), Jiang et al. (2024) and industry insights. Inspired by these requirements, we design and implement OR-Instruct, a semi-automated process for creating synthetic data tailored to optimization modeling. The process uses an iterative bootstrapping algorithm (see Figure 2). Initially, we collect a set of seed industry cases (e.g., 686 cases in our study) and add them to the training data pool. Following this, we use two strategies. One is expansion, which employs GPT-4 to generate data covering a wider range of scenarios and question types, expanding to thousands of scenarios and tens of question types. While this strategy broadens the coverage of scenarios, it does not ensure diversity in problem difficulty. GPT-4 struggles with independently solving complex problems and tends to simplify the optimization model, leading the strategy to generate results predominantly for easy and medium-level problems. To ensure that synthesized data retains the difficulty of the original seed data, we introduce the augmentation strategy. This strategy involves modifying objectives and constraints, rephrasing questions, and incorporating diverse modeling techniques. Essentially, we decompose the remaining requirements into subtasks, enabling GPT-4 to generate examples with greater problem-solution diversity within its capabilities. Finally, heuristics are used to automatically filter out obviously low-quality data. This cycle can be repeated multiple iterations until the training data pool reaches the requisite volume (e.g., 32,481 cases in our study). Importantly, this process is designed to facilitate customization by simply introducing new seed data and scenarios as needed.
    Figure 2: Overview of OR-Instruct.
  2. 2.
    To evaluate the efficacy of OR-Instruct, we introduce IndustryOR, the inaugural industrial benchmark tailored for optimization modeling. This benchmark uses data sourced from 13 different industries, covering 5 types of questions across 3 levels of difficulty. For comprehensive evaluation, we also include NL4OPT (Ramamonjison et al. 2023) and MAMO (Huang et al. 2024b) benchmarks.
  3. 3.
    We utilize the synthetic data generated by OR-Instruct to train a series of open-source large language models, each approximately 7 billion parameters in size, including Mistral-7B (Jiang et al. 2023), Deepseek-Math-7B-Base (Shao et al. 2024), and LLaMA-3-8B (AI@Meta 2024). We designate these trained models as Operation Research Language Models (ORLMs) and observe a marked enhancement in their optimization modeling capabilities. Our experimental results show that our best-performing ORLM achieves state-of-the-art performance on the NL4OPT, MAMO, and IndustryOR benchmarks.
To our best knowledge, our work is the first initiative to train open-source LLMs specifically for optimization modeling in real-world industries. This approach significantly reduces the reliance on closed-source LLMs thus adding significant flexibility to the training process and greatly enhancing privacy preservation for industrial applications. Remarkably, our highest-performing ORLM achieves state-of-the-art results on the NL4OPT, MAMO, and IndustryOR benchmarks. This indicates that even models with a parameter count as moderate as 7 billion can surpass the capabilities of larger models such as GPT-4 and comparable agent-based frameworks, following our specialized training methodology. This underscores the effectiveness of our approach.
We also want to comment that our proposed training method, the OR-Instruct, has the ability to facilitate customization to address particular needs of larger models, such as supply chain management, scheduling, inventory issues, or enhancing capabilities in linear programming, integer programming, and mixed integer programming. This synthetic data framework robustly addresses the scarcity of training data for optimization modeling, laying a solid foundation for future advancements in deep learning model training in this field.
Finally, we note that almost all of our results, including certain samples from the training data set, industry-level benchmark for optimization modeling, and all the best-performing LLMs that have been fine-tuned, are made open source. We hope that our work can serve as a foundation for subsequent explorations, empowering future researchers to enhance the utility and efficiency of large language models in optimization modeling.
The structure of this paper is organized as follows: Section 2 reviews the literature pertinent to our research. Section 3 delineates four essential conditions required for the training data and outlines the implementation specifics of the semi-automated synthetic data framework for optimized modeling, OR-Instruct. Section 4 examines the impact of training open-source large models on various test sets and presents the experimental outcomes of customized enhancements in modeling capabilities. Section 5 compares the modeling effectiveness of GPT-4 with our ORLM and validates the necessity of the four key conditions through ablation studies. We conclude the paper in Section 6.
2 Literature Review
Our research is related to several streams of literature, we will provide a detailed review of each stream below.
AI approach for facilitating solution to operations research problems. With the fast development of AI techniques, there have been a growing amount of research leveraging advanced techniques, such as deep learning, to address challenging operational research (OR) problems traditionally resistant to conventional methods. For example, in the supply chain context, Gijsbrechts et al. (2022) employ deep reinforcement learning to enhance inventory management, demonstrating parity with leading heuristics and approximate dynamic programming approaches. Han et al. (2023) examine a data-driven newsvendor problem resolved through deep neural networks (DNNs), offering theoretical assurances on excess risk bounds and validating the method’s effectiveness, Qi et al. (2023) investigate a data-driven multiperiod inventory replenishment problem with uncertain demand and vendor lead time, propose a one-step end-to-end framework that uses deep learning models to output the suggested replenishment amount directly from input features. In the revenue management field, Aouad and Désir (2022) propose a neural network-based discrete choice model called RUMnets, which combines the theoretical robustness of traditional economic models with the flexibility and learning capabilities of modern artificial intelligence approaches, leading to better predictions and insights into consumer behavior. Wang et al. (2023a) explore the deployment of deep neural networks in making the optimal assortment selection, demonstrating its effectiveness in capturing consumer choice. As for queuing theory, Dai and Gluzman (2022) investigate the deployment of advanced policy gradient methods for controlling multi-class queueing networks, underscoring the superior efficacy of the PPO algorithm under diverse load scenarios. For personalized recommendation, Wang et al. (2023c) utilize a novel unsupervised deep learning network and graph embedding techniques to improve recommendation accuracy and diversity by effectively transferring check-in patterns across different geographical contexts. Regarding general Markov decision process, Zhang et al. (2021) introduce a model-free reinforcement learning algorithm that attains asymptotically optimal sample complexity for learning an $\epsilon$-optimal policy within discounted Markov decision processes (MDPs). In addition, these techniques have been used in demand prediction, for example, Lee et al. (2024) propose a Transformer-based methodology for time series conformal prediction, utilizing its decoder to estimate prediction intervals through the quantile prediction of residuals, showcasing enhanced performance over existing conformal prediction techniques. Chen et al. (2020) introduce a novel architecture of convolutional neural networks (CNNs) that significantly improves forecasting by modeling the conditional distribution of future series data directly, rather than relying on traditional autoregressive methods. AI technology also demonstrates significant strengths in facilitating the implementation of operations research, for instance, Liu et al. (2023) employ reinforcement learning combined with spatial data to optimize the allocation of resources and strategize for disease outbreak responses. Parmentier (2022) introduces a novel paradigm that integrates machine learning with operations research to efficiently approximate complex industrial problems by transforming them into classic, more tractable optimization challenges.
In addition, there have also been recent attempts in using AI methods to empower optimization algorithm design. For instance, deep learning or machine learning can be used to accelerate conventional optimization algorithms for tackling mixed-integer programming problems (e.g., Chen et al. 2023b, Nair et al. 2020). Numerous researchers also explore end-to-end optimization problem-solving strategies based on deep learning methods, motivated by the high structural similarity among many optimization problems, with differences primarily in data distribution (e.g., Maragno et al. 2023, Chen et al. 2022a, Khalil et al. 2017). Compared to this work, the above stream of works studies the integration of AI approach and traditional OR approaches in solving OR problems. In contrast, the main focus of the present work is in the facilitation and automation of the modeling process of the optimization model, which is different from the above stream.
Automated optimization modeling is an emerging area within the field of Decision Intelligence, which aims to automate mathematical modeling and solving tasks to facilitate efficient decision-making. Recently, a plethora of studies have emerged, aiming to bridge the gap between natural language processing and mathematical optimization. One of the pioneering endeavors in this field is the NL4OPT competition, which focuses on formulating mathematical models for optimization problems (Ramamonjison et al. 2023). In particular, their work introduces a two-step framework using PLMs and offers a widely used benchmark for intelligence mathematical modeling. Expanding on this, Xiao et al. (2023) introduce Chain-of-Experts, a multi-agent LLM framework to build optimization model that significantly outperforms GPT-4 by incorporating domain-specific knowledge and cooperative reasoning strategies. AhmadiTeshnizi et al. (2024) propose OptiMUS, a robust LLM-based agent designed to formulate and solve MILP problems. This system not only develops mathematical models but also assists in writing and debugging solver code, showcasing a comprehensive solution that extends beyond simple problem formulation. In a similar vein, Li et al. (2023a) develop a framework called OptiGuide which leverages LLMs to interpret and explain optimization outcomes in supply chain management, thereby enhancing stakeholder comprehension and trust. In terms of diagnosing and resolving optimization issues, Chen et al. (2023a) introduce OptiChat, a natural language-based system equipped with GPT-4, which assists in identifying and addressing infeasible optimization models. This tool significantly aids both experts and novices in understanding and refining their models. Unlike these methods, our approach involves training open-source LLMs and delivering a complete solution using direct prompting, which is more concise and privacy-preserving than multi-agents interaction and simply calling API from advanced LLMs like GPT-4.
Synthetic Data is one of the most important methods to improve the performance of large language models, with the advantage of low cost and diversity. A review of literature on synthetic data and LLMs is referred to Ding et al. (2024). In the realm of synthetic data for optimization modeling, traditional approaches often necessitate specialized expertise to transform real-world issues into mathematical formulations, thereby constraining broader accessibility and adaptability. Predominantly, researches explore how to automatically synthesize mixed-integer linear programming (MILP) to address these challenges effectively. For example, the methods proposed by Pawlak and Krawiec (2017) and Sroka and Pawlak (2018) utilize MILP examples for the synthesis of constraints. These methods develop an approach that can generate constraints from examples of both feasible and infeasible solutions, thereby demonstrating the feasibility of automating this conversion process. Further expanding on this theme, Pawlak and O’Neill (2021) leverage evolutionary strategies to generate constraints that improve the adaptability and efficiency of optimization models, thus enhancing the model’s flexibility in handling complex scenarios. Prasath and Karande (2023) examine the effectiveness of synthesizing mathematical programs from natural language specifications based on NL4OPT competition dataset, focusing on data synthesis methods using the CodeT5 model with enhancements from data augmentation and post-processing. The study highlights how synthetic data generation, notably through techniques like reverse translation and parameter variation, significantly improves model performance in solving optimization problems. The proposed synthetic data framework in Li et al. (2023b) leverages LLMs to identify decision variables, classify constraints, and generate MILP models based on NL4OPT problems. This approach significantly outperforms traditional methods, achieving higher accuracy in constraint recognition and model formulation. In contrast to the aforementioned synthetic data approaches, the method proposed in this paper is designed for the general purpose of synthesizing high quality data for optimization modeling from unstructured natural language descriptions. This allows for the development of synthesis strategies that are not only tailored to specific problem but also capable of customized data augmentation.
3 Semi-Automated Framework for Data Generation
In this section, we present the formal definition of the optimization modeling task. We first discuss some critical requirements for effectively training an open-source LLM to perform this task. Based on these requirements, we design a semi-automated data synthesis framework OR-Instruct which achieves the desired properties.
3.1 Data Requirements
An Expected Training Example Input - Natural Language Problem: A company has three transportation options to choose from to transport $25$ tons of cargo, namely trucks, airplanes, and ships with costs \$$\$100,\$120,\$130$ per ton and capacities of $10,20,30$ tons respectively. The company can’t choose trucks and ships together. How should the company optimize the selection and allocation of these methods to minimize overall costs? Target - Mathematical Model and Program: ## Mathematical Model:
### Variables:
- $x_{1},x_{2},x_{3}$: 0-1 variables indicating whether trucks, airplanes, and ships are selected, respectively.
- $y_{1},y_{2},y_{3}$: Non-negative continuous variables indicating the volume of cargo transported by three modes.
### Objectives and Constraints:
[…]
Figure 3: An expected training example for optimization modeling task.
Given an OR problem $p$ described in natural language, optimization modeling (Berry and Houston 1995, AhmadiTeshnizi et al. 2024) involves constructing a mathematical model $m$ that converts the real-world problem into an optimization model with concrete decision variables, objective function and constraints. Then, for large-scale problems, in order to obtain an optimal solution, we further need to convert the mathematical model into a program $c$ compatible with a specific optimization solver. Hence, an expected training example for this task is usually required in the form of the triplet $(p,m,c)$, as illustrated in Figure 3. Training a large language model $f$ for this task essentially involves learning a function $f:p\rightarrow(m,c)$ that maps problems to their corresponding mathematical models and solver programs. In this paper, without loss of generality, COPT is employed as the default solver. COPT (Ge et al. 2022), an acronym for Cardinal Optimizer, is able to tackle large-scale optimization problems and achieves good performance in a variety of tasks (Hans 2002). We select the COPT solver because the predominant large language models available in the market have not yet mastered the COPT syntax. This choice underscores the efficiency and extensibility of our framework. However, any well-established solver may be utilized in the training dataset construction.
In addition to the desired format as described above, a desired training dataset should demonstrate effectiveness, robustness, and practical applicability to real-world scenarios. In particular,
Requirement 1
Comprehensive Coverage: The dataset should cover: 1) diverse scenarios such as supply chain optimization, scheduling, inventory management, and transportation logistics; 2) different problem types like linear programming, integer programming, and mixed integer programming; and 3) varying difficulty levels (easy, medium, hard) — we provide a way to describe the difficulty level of a problem in Appendix 8.1.
Requirement 2
Environmental Adaptability: In real-world industrial settings, the objectives and constraints of problems often change due to shifts in business goals, market conditions, or resource availability. A good optimization model should be able to adapt to those changes easily. Therefore, it is vital that the dataset includes cases reflecting these dynamic changes.
Requirement 3
Linguistic Diversity: Problems described in natural language often show different syntax, ambiguities, and complexities. For example, one problem might mention "inventory overflow" while another refers to "excess stock". Including this linguistic diversity in the dataset could improve the model’s ability to understand varied descriptions.
Requirement 4
Technique Variability: For some challenging problems, there may be multiple modeling techniques, such as linearizing a nonlinear problem by introducing auxiliary variables, or introducing the big $M$ term in an integer optimization formulation. Including this variety in the dataset allows the model to learn different modeling techniques and approaches.
3.2 OR-Instruct: Towards Training Effective OR LLMs
In the above section, we describe some critical requirements on the dataset for an effective training of an LLM to be equipped with good optimization modeling capabilities. However, collecting data on a large scale that aligns with the above requirements presents significant challenges because: (1) such data primarily resides within private industrial cases, and no public datasets sufficiently meet the requirements, (2) training data suitable for optimization modeling is limited and requires extensive time to collect, and (3) conventional synthesis methods (for example, Wang et al. 2022) encounter substantial obstacles in generating this type of data, as will be elaborated below. In the following, to address the above challenges, we design and implement OR-Instruct, a semi-automated process for creating synthetic data tailored to these requirements. The overall pipeline of OR-Instruct is depicted in Figure 2, which iteratively applies two strategies, namely expansion and augmentation, to the training data pool, followed by the filtering of obviously low-quality data.
Strategy 1: Expansion for Comprehensive Coverage.
Initially, OR-Instruct attempts to generate new data by expanding scenarios and question types from the training data pool in a bootstrapping fashion with GPT-4. In our particular case, we start with 686 real-world industry cases collected from some OR textbooks and our previous industrial projects (after proper abstraction and anonymization). For each generation, we sample 3 entries from this pool as in-context examples, keeping the input token length suitable for GPT-4. Of the 3 entries, 2 are from real-world entries and 1 is from the model-generated entries in previous iterations, if available, to promote diversity. The prompting template is shown in Appendix 8.3.
This approach is similar to general synthesis methods such as Wang et al. (2022), but OR-Instruct focuses on expanding scenarios rather than tasks. We observe that this step partially addresses the Comprehensive Coverage requirement in terms of scenarios and question types. However, it falls short of meeting the other requirements, especially concerning varying levels of difficulty. In manually reviewing the difficulty of generated example, of 50 cases, 87% are deemed easy, 13% medium, and none hard, as judged by criteria in Appendix 8.1. We also provide examples in Appendix 8.9 for comparing generated easy entries with real-world hard entries. Fortunately, the original seed data already shows a diverse range of difficulties. Thus, we can naturally enhance the difficulty diversity by augmenting them described in the next strategy, thereby further addressing Comprehensive Coverage.
Strategy 2: Augmentation for Problem-Solution Diversity.
To fulfill the remaining requirements, the OR-Instruct adopts an augmentation strategy for the original seed data, designed to both enhance the diversity and robustness of the dataset while preserving the complexity inherent in optimization modeling problems. This strategy is crafted to mirror potential scenarios that might be encountered in real-world industrial applications, encompassing modifications such as alterations in problem descriptions, model adjustments, or concurrent changes. To address these three scenarios, the implementation of this strategy involves three specific tactics: rephrasing questions, modifying objectives and constraints, and integrating diverse modeling techniques. Together, these tactics aim to broaden the range of problem-solution diversity, thereby enriching the training dataset.
  1. (a)
    Altering Objectives and Constraints: The first augmentation involves adding, removing, or replacing objectives and constraints in the problem, along with making necessary adjustments to the mathematical models and programs. Specifically, we start by providing GPT-4 with the original example and ask it to list five potential changes to the objectives and constraints. These suggested changes are then fed back to GPT-4 using a prepared few-shot prompt to modify the problem, model, and programs accordingly. This augmentation is designed to enhance Environmental Adaptability. An example is provided in Figure 4, and the corresponding prompting template is shown in Appendix 8.4.
    Altering Objectives and Constraints for Requirement 2 Original: Q: … The company can’t choose trucks and ships together. Denote the cost … Augmented: Q: … The company can’t choose trucks and ships together. Due to the special nature of the goods, the company has decided that if trucks are chosen, airplanes must also be selected for transportation. Denote the cost … A: … New dependency constraint (choosing trucks necessitates choosing airplanes): $x_{1}\leq x_{2}$ 1... 2model.addConstr(x[’trucks’] <= x[’airplanes’], name="New constraint") 3...
    Figure 4: An example illustrating altering objectives and constraints. Italicized text denotes altered content.
  2. (b)
    Rephrasing Questions: The second augmentation modifies the description of the question, simulating the expression habits of different users. This process involves instructing GPT-4 to rewrite the target problem, either simplifying or complicating it, while ensuring the core logic aligns with the solution, including the mathematical model and programs. This augmentation is designed to enhance Linguistic Diversity. An example is provided in Figure 5, and the corresponding prompting template is shown in Appendix 8.5.
    Rephrasing Questions for Requirement 3 Original: Q: A company has three transportation options to choose from to transport $25$ tons of cargo, namely trucks, airplanes, and ships with costs \$$\$100,\$120,\$130$ per ton and capacities of $10,20,30$ tons respectively. The company can’t choose trucks and ships together. How should the company optimize the selection and allocation of these methods to minimize overall costs? Augmented: Q: A corporation wants to transport 25 tons of cargo with least cost, and must choose from three transportation modes: trucks, airplanes, and ships. These options cost \$100, \$120, and \$130 per ton, respectively, with capacities of 10, 20, and 30 tons. However, trucks and ships cannot be used together.
    Figure 5: An example illustrating rephrasing questions. Italicized text denotes the question that has been rephrased.
  3. (c)
    Incorporating Multiple Modeling Techniques: The third augmentation explores the use of different modeling techniques. We identify five potential techniques from engineers’ experiences, such as introducing auxiliary variables or using the Big $M$ method, for GPT-4 to choose from in modifying an objective or constraint in the original mathematical model. This augmentation is designed to enhance Solution Variability. An example is provided in Figure 6, and the corresponding prompting template is shown in Appendix 8.6.
    Incorporating Multiple Modeling Techniques for Requirement 4 Original: A: Mutual exclusion constraint (trucks and ships cannot be selected simultaneously): $x_{1}+x_{3}\leq 1$ Augmented: A: Mutual exclusion constraint (Using big $M$ method): $x_{1}\leq(1-x_{3})M,\text{where }M\text{ is a large number}$ 1... 2model.addConstr(x[’trucks’] <= (1-x[’ships’])*M, name="New constraint") 3...
    Figure 6: An example illustrating incorporating multiple modeling techniques. Italicized text denotes the section that have been re-modeled using new techniques.
Postprocessing and Filtering
Finally, OR-Instruct implements several heuristic measures on the generated examples. Initially, we eliminate examples that contain duplicate questions or those that overlap with questions from our evaluation benchmarks. Subsequently, we apply a regular match correction function to amend minor grammatical errors in the programs, which are occasionally induced by GPT-4’s limited familiarity with the COPT API (refer to Appendix 8.2). Training data that contain unexecutable programs are discarded, as they represent manifestly low-quality data. After conducting a manual review of the remaining data, which demonstrates an acceptable accuracy rate — 70% for expansion data and 75% for augmentation data — we opt to bypass further filtering in pursuit of establishing a fully automatic filtering mechanism. This final filtering step is projected to exclude approximately 39% of the generated examples, with the surviving examples subsequently incorporated into our training data pool.
4 Results
In this section, we train the open-source 7b size large language models with synthetic data from OR-Instruct and compare the results with GPT-4 as well as benchmarks. Then, using the case of mixed integer linear programming as an example, we demonstrate the customized enhancement effect of OR-Instruct by including directionally generated training data.
4.1 Data Generation by Using OR-Instruct
In our experiment, we initiate the study with 686 cases derived from previous industry projects, which the team engages with previously. These cases are abstracted and anonymized appropriately. Throughout the OR-Instruct process, we utilize the proprietary LLM gpt-4-0613 as the standard model. For each cycle, the training data pool undergoes an expansion procedure 20,000 times and each augmentation operation is applied 6,000 times. Subsequently, an automatic filtration system excludes low-quality entries. After completing two iterations of this framework, we amass a total of 32,481 training examples.
[Distribution of industies] Refer to caption [Question type] Refer to caption
Figure 7: Statistics of the data generated by OR-Instruct
Statistics: Figure 7 presents the statistics of the OR-Instruct Data. Of the data, 57% is generated by the expansion operation, 17.2% by the augmentation of altering objectives and constraints, 15.3% by the augmentation of rephrasing questions, and 10.5% by the augmentation of incorporating multiple modeling techniques. For scenario diversity, we have abstracted 16 industries from a total of 1,556 expanded scenarios to facilitate their representation in a plot. Regarding question type diversity, it expands to 8 question types. In terms of data quality, the accuracy of correctness are 70% for expansion data and 75% for augmentation data. We find these accuracy acceptable in exchange for automatic filtering and more cost-effective data synthesis. Additionally, our preliminary experiments have demonstrated the benefits of incorporating such data.
4.2 Model Training and Inference
We borrow a widely used training framework from open-instruct (Wang et al. 2023d.) where we take a natural language OR problem wrapped in an Alpaca-like template (Taori et al. 2023.) as the input prompt (see Appendix 8.7), and treat a complete solution including mathematical models and programs as the target completion. During training, we compute the loss only over the target completion. We apply OR-Instruct Data to train several open-source LLMs of approximately 7b size, including Mistral-7B (Jiang et al. 2023.), Deepseek-Math-7B-Base (Shao et al. 2024.), and LLaMA-3-8B (AI@Meta 2024.), as the backbones. We list the hyper-parameters for each backbone in Table 1. We call the resulting models as ORLMs. For inference during evaluation, we always use greedy decoding under 0-shot setting to eliminate randomness, selecting the top-1 completion as the final solution. We extract the program from the solution and execute it to obtain the predicted optimal value.
Table 1: Hyper-parameters for training ORLMs.
Backbone BatchSize LearningRate Epochs
Mistral-7B 512 3e-6 2
Deepseek-Math-7B-Base 128 2e-5 2
LLaMA-3-8B 64 5e-6 2
4.3 Evaluation and Baselines
Evaluation Benchmarks and Metrics: We use NL4OPT (Ramamonjison et al. 2023.), MAMO (Huang et al. 2024b.), and IndustryOR as evaluation benchmarks. NL4OPT is the most widely used benchmark for operations research and includes 289 linear programming problems in its test set. However, NL4OPT only provides mathematical models as targets, which complicates the verification of execution accuracy due to the absence of optimal solutions. To address this, we convert these mathematical models into programs using GPT-4, calculate and check the optimal solutions, and use these as ground truth. MAMO, a concurrent project, evaluates the mathematical modeling capabilities in LLMs. It includes 652 easy and 211 complex linear programming problems, each paired with its corresponding optimal solution, sourced from various academic materials. IndustryOR, which is proposed in this work, is the first industrial benchmark, consisting of 100 real-world OR problems from eight industries. It covers 5 types of questions linear programming, integer programming, mixed integer programming, non-linear programming, and others across 3 levels of difficulty. We measure performance using execution accuracy, where an executed optimal value that matches any provided ground truth optimal value is considered correct. Compared to NL4OPT, this metric enables a fully automated evaluation and provides greater flexibility for mathematical modeling approaches.
Baselines: To ensure a comprehensive evaluation, we select a diverse set of models from previous methods for comparison. We include tag-BART (Kani and Gangwar 2022), which secured the first place in the NeurIPS competition (Ramamonjison et al. 2022). Additionally, we consider methods that utilize proprietary LLMs. The "Standard" prompting method involves prompting a proprietary LLM to produce mathematical programs, serving as a fundamental baseline. We also incorporate complex prompt engineering methods such as Reflexion (Shinn et al. 2023), Chain-of-Experts (Xiao et al. 2023), and OptiMUS (AhmadiTeshnizi et al. 2024). These methods employ agents to refine both mathematical models and programs, achieving good performances on NL4OPT. We report their performance based on GPT-3.5 and GPT-4, respectively. Note that we implement the standard prompting on IndustryOR using the toolkit released by Chain-of-Experts (Xiao et al. 2023).
Table 2: Comparison of performance on the NL4OPT, MAMO, and IndustryOR benchmarks. Values marked with a * are directly copied from original papers, with blanks where data were not reported. The highest results are highlighted in bold.
Method NL4OPT MAMO EasyLP MAMO ComplexLP IndustryOR Micro Avg Macro Avg
Methods based on PLMs
tag-BART 47.9%* - - - - -
Methods based on GPT-3.5
Standard 42.4%* - - - - -
Reflexion 50.7%* - - - - -
Chain-of-Experts 58.9%* - - - - -
Methods based on GPT-4
Standard 47.3%* 66.5%* 14.6%* 28.0% 50.2% 39.1%
Reflexion 53.0%* - - - - -
Chain-of-Experts 64.2%* - - - - -
OptiMUS 78.8%* - - - - -
ORLMs based on open-source LLMs
ORLM-Mistral-7B 84.4% 81.4% 32.0% 27.0% 68.8% 56.2%
ORLM-Deepseek-Math-7B-Base 86.5% 82.2% 37.9% 33.0% 71.2% 59.9%
ORLM-LLaMA-3-8B 85.7% 82.3% 37.4% 38.0% 71.4% 60.8%
The results are presented in Table 2. First, it is clear that methods based on LLMs generally outperform the PLM-based best method (tag-BART) in the NL4OPT test. This suggests that PLMs have limited generalization capabilities. For proprietary LLMs, as the mathematical reasoning capability increases from GPT-3.5 to GPT-4, we observe that performance has advanced across all prompt engineering methods. Finally, ORLMs based on various open-source LLMs have demonstrated significantly improved optimization modeling capabilities, compared to vanilla open-source LLMs, which score 0 for all benchmarks in our preliminary experiments due to failing to output executable programs. This underscores the effectiveness of the OR-Instruct data. Our best-performing ORLM, trained on LLaMA-3-8B, achieves state-of-the-art performance on NL4OPT, MAMO (including both easy and complex linear programming), and IndustryOR benchmarks, surpassing Standard prompting based on GPT-4 by 42.2% in micro average [^1] Micro average aggregates the outcomes of all classes to compute the metrics, heavily weighing the classes with more instances. This method is effective for evaluating the overall effectiveness of the model, particularly in datasets where some classes significantly outnumber others. and 55.4% in macro average [^2] Macro average calculates individual metrics for each class without regard to class frequency, and then averages these metrics, treating all classes with equal importance. This approach highlights model performance on minority classes and is beneficial for assessing model fairness across diverse class distributions.. We also want to highlight that, unlike tag-BART, we ensure no training and validation data from NL4OPT was used. Thus, the performances of the ORLMs across all benchmarks reflect out-of-domain generalization.
4.4 Customization Enhancement
In this subsection, we illustrate the capability of our framework to enhance large models in specific areas by generating targeted types of optimization modeling data. Specifically, we focus on the scenario of mixed integer linear programming (MILP). We provide an empirical example demonstrating how the quality of responses improves with training. Initially, we use a base dataset comprising 10,000 general modeling cases. To enhance the ORLM’s proficiency in MILP, we selected 10 fundamental MILP questions and their answers for inclusion in the seed data pool. After several iterative rounds, this process yields 2,000 tailored training data points for MILP. We will subsequently assess the impact of incorporating these 2,000 data points into the training dataset by comparing the ORLM’s performance before and after their addition.
Figure 8: Comparison of ORLM before and after adding specific data for training and GPT-4 responses, the code part is omitted here. Incorrect responses are marked in red.
As depicted in Figure 8, in response to the scenario presented in Figure 1, GPT-4 inaccurately interprets logical relationships as quantitative ones and includes nonlinear terms during the modeling process as shown in the red area, adversely affecting solver performance. A similar mistake is observed in the general ORLM, which also misinterprets logical relationships as quantitative. In contrast, the targeted-trained ORLM accurately transforms the problem into a correct mathematical model by utilizing indicator variables and the Big $M$ method. This showcases the robust generalization capability of OR-Instruct. Furthermore, OR-Instruct not only facilitates customized improvements across different modeling types but also demonstrates potential for targeted enhancements within specific industrial sectors.
5 Analysis and Discussion
5.1 Detailed Comparison of ORLM vs GPT-4 on IndustryOR
To assess the optimization modeling capabilities across different levels of difficulty and question types, we compare ORLM-LLaMA-3-8B and Standard-GPT-4 on IndustryOR as shown in Figure 9. ORLM-LLaMA-3-8B shows superior performance over Standard-GPT-4 across all difficulty levels, especially in the hard category. Regarding different question types, ORLM-LLaMA-3-8B outperforms Standard-GPT-4 in linear programming, integer programming, and mixed-integer programming. Both models perform poorly in non-linear programming and other rare question types. This is partly due to the fact that the seed dataset collected contains too few examples of nonlinear and other types of optimization modeling, resulting in a low percentage of that type in the training dataset, and partly due to the fact that the inherent complexity and scarcity of these types of questions. Overall, the OR-Instruct data proves effective in enhancing Comprehensive Coverage across various question types and difficulty levels.
Figure 9: Comparison of ORLM and GPT-4 on IndustryOR across different difficulty levels and question types.
5.2 Ablation Study on OR-Instruct Augmentations
To further verify the effectiveness of the augmentations in OR-Instruct, we conduct detailed ablation experiments to study the effects of altering objectives and constraints, rephrasing questions, and incorporating multiple modeling techniques. Specifically, we select a controlled sample size of 3,000 with certain augmentations from the OR-Instruct data, and then use these to train the LLaMA-3-8B model with the same hyper-parameters across all experiments. The results are presented in Table 3. Training data with all three augmentations (denoted as Full Augmentations) achieves a base performance of 68.6% in micro average and 55.7% in macro average. Removing any of the three augmentations from the base setting leads to a performance drop across all benchmarks, both in micro and macro averages. Rephrasing questions seems slightly more important than the other two. Overall, the results show that all three augmentations contribute to general performance, proving their effectiveness in enhancing the data diversity.
Table 3: Ablation study on OR-Instruct Augmentations.
Method NL4OPT MAMO EasyLP MAMO ComplexLP IndustryOR Micro Avg Macro Avg
Full Augmentations 78.3% 80.6% 43.1% 21.0% 68.6% 55.7%
w/o Altering Obj&Const 77.5% 79.2% 36.4% 20.0% 66.4% 53.2%
w/o Rephrasing Questions 74.2% 77.3% 41.1% 15.0% 65.1% 51.9%
w/o Multiple Modeling 78.3% 78.0% 38.8% 18.0% 66.2% 53.2%
6 Future Direction
Large language models (LLMs) are increasingly influential across various industries, and the rise of privatized and domain-specific LLMs presents new topics for management research. This work represents the inaugural attempt to train open-source large language models to the domain of optimization modeling, liberating the field from reliance on proprietary advanced models, which facilitates the application and commercialization of large models within industrial sectors. We hope this work shed lights on the deep integration of large language models with operations research and management science, potential future research directions are listed as follows:
  • Incorporating Sophisticated Techniques. Building upon our foundational work in addressing the scarcity of training data, future research could delve into more advanced techniques. These might include training multiple specialized large language models (LLMs) for collaborative multi-agent modeling, refining transformer architectures to enhance reasoning abilities, and utilizing alignment algorithms like reinforcement learning from human feedback (RLHF), direct preference optimization (DPO), and Kahneman-Tversky optimization (KTO) to boost the performance of large models. Such investigations could unlock the potential of large models to surpass traditional methods in tackling complex operations research issues. Additionally, these strategies could be adapted to address emerging industrial challenges, such as predictive tasks in machine learning.
  • Knowledge Management and Human-Machine Collaboration. Large language models (LLMs) exhibit superior learning efficiency and expansive knowledge storage, heralding a new era in human-computer collaboration and enterprise management. LLMs can leverage technologies such as retrieval-augmented generation (RAG) for historical knowledge management, effectively preserving corporate memory and mitigating the impact of employee turnover. Their advanced text mining capabilities have the potential to revolutionize market research analysis and report generation. This leads to human-machine interactions that transcend the knowledge encapsulated within trained LLMs by integrating a company’s accumulated historical experiences and external human-derived information. Such integration facilitates more efficient task execution, such as optimization modeling, thereby driving digital transformation and fostering innovation within enterprises.
  • Consumer Behavior Insights through LLM. Large models’ capability to simulate human social behaviors offers valuable insights for enterprises analyzing consumer behavior and formulating market strategies. For instance, through simulations, large language models can generate preferences that closely resemble human behaviors, assisting businesses in conducting comprehensive A/B testing. Additionally, the inherent choice models in algorithms such as RLHF facilitate the capture of consumer choice behaviors, providing new perspectives on revenue management issues, including product assortment.
  • Economic and Game Theoretical Aspects of LLM. The commercialization of large models introduces new research dimensions in management studies. The emergence of large models transforms specific industry textual data, such as project reports, into a valuable new data asset. Managing, pricing, and transacting these data assets remain unresolved challenges. Additionally, the advent of large models creates new game relationships between R&D companies and service companies, employees and LLMs, as well as LLMs and customers, related problems in this field like contract theory, mechanism design, and game theory are still open.
7 Conclusion
In this paper, we propose a new path for the training of open-source large language models (LLMs) specifically tailored for optimization modeling. We delineate four critical requirements essential for the training dataset of optimization modeling LLMs and develop OR-Instruct, a semi-automated framework for generating synthetic data that meets these specific needs. Additionally, we introduce the IndustryOR benchmark, the first of its kind in the industry. We utilize the OR-Instruct data to train open-source LLMs with approximately 7 billion parameters, which significantly enhances their optimization modeling capabilities, achieving state-of-the-art performance across all test datasets.
This research addresses the issue of insufficient training data in optimization modeling. Looking ahead, we plan to explore potential enhancements through diverse architectures, agent-based models, and advanced modeling techniques to elevate modeling capabilities, leveraging the training dataset provided by OR-Instruct. Additionally, we will examine managerial innovations and practical challenges arising from the real-world application of large language models in industrial contexts.
Acknowledgement
This research is partially supported by National Natural Science Foundation of China (Grants 72394360, 72394364, 72171141), the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing. Zizhuo Wang is partially supported by the National Natural Science Foundation of China (NSFC) [Grant NSFC-72150002, NSFC-72394361] and the Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (2023B1212010001). Dongdong Ge is partially supported by the National Natural Science Foundation of China (NSFC) [Grant NSFC-72150001, 72225009, 72394360, 72394365].